Celia and Dina start at the same time from two towns 90km apart, travelling
towards each other. Celia’s rate is 5km faster than Dina’s rate. They met
after two hours. Find the average speed of each. Please include Represent, Relate, Equate, Solve and Prove Thank you!

Answer by ankor@dixie-net.com(22740)

Celia and Dina start at the same time from two towns 90km apart, travelling
towards each other.
Celia’s rate is 5km faster than Dina’s rate.
They met after two hours
:
let s = Dina's speed
Celia goes 5 km/hr faster, therefore
(s+5) = Celia's speed
:
When they meet their total dist is 90 km. Write a distance equation
dist = time * speed
2s + 2(s+5) = 90
2s + 2s + 10 = 90
4s = 90 - 10
4s = 80
s = 80/4
s = 20 km/hr is D's speed
then
20 + 5 = 25 km/hr is C's speed
:
:
Check this, find the actual dist each traveled
2 * 20 = 40 km
2 * 25 = 50
------------
total d: 90 km